OpenStudy (kayders1997):

Please help will fan and medal!!! The altitude of a triangle is increasing at a rate of 1 cm/min while the area is increasing at a rate of 2 cm^2/min. At what rate is the base of the triangle changing when the altitude is 10 cm and the area is 100cm^2 okay for this one I made a sketch I had a triangle and the 10 cm on the y side 1 cm/min there too and than the 100 cm^2 in the middle with the 2cm^2/min and than I used the area formula and differentiated A=1/2bh

OpenStudy (danjs):

yeah sketch it, and list all the given things , rates, and values

OpenStudy (kayders1997):

Okay

OpenStudy (danjs):

da/dt = 1 cm/min dA/dt = 2 cm^2/min db/dt = ?? when a=10 cm, and A=100cm^2 -------------------------------- Given the rates of change for the altitude and the area. Find the rate of change of the base, at the instant the altitude is 10 and the area is 100

OpenStudy (kayders1997):

I wish I could send a picture of my sketch

OpenStudy (danjs):

i can picture it, the altitude is getting larger, and in tern the area also increases

OpenStudy (kayders1997):

Right

OpenStudy (kayders1997):

So to find side b the base can you do...? Area formula? 1/2bh=100cm^2 when h=10?

OpenStudy (kayders1997):

And would you do that formula for the first part? Differentiation for the area equation?

OpenStudy (kayders1997):

I'm not very good at these, its a pretty concept for us

OpenStudy (danjs):

the general way to solve these related rate probs, is to understand what you have and what to find Here the Area function relates all three things , area , altitude

OpenStudy (danjs):

you need a relationship between the different variabes, in this prob, we need the derivative w.r.t. time d/dt of the total area function, none of the terms are in terms of a time, base, height, area have to use the chain rule to get to the time derivative

OpenStudy (kayders1997):

Okay

OpenStudy (kayders1997):

I used A=1/2bh to get da/dt=1/2(b(dh/dt)+h(db/bt))

OpenStudy (kayders1997):

Idk if that is right

OpenStudy (danjs):

d/dt [A] = d/dt [ b * h] yah looks good

OpenStudy (kayders1997):

Yay :)

OpenStudy (kayders1997):

Than I multiplied that 1/2 through

OpenStudy (kayders1997):

So when I started plugging in numbers I got 2=1/2b(dh/dt)+q/2(10)(1)

OpenStudy (kayders1997):

1/2 after the plus sign

OpenStudy (kayders1997):

And now would I use A=1/2bh when h=10 and A=100 to find b?

OpenStudy (danjs):

yeah you have \[\frac{ dA }{ dt }= \frac{ 1 }{ 2 }*[b*\frac{ dh }{ dt } + h*\frac{ db }{ dt }]\] you know the height and the area at a moment in time, A=1/2 b * h 100 = 1/2 * b * 10 so the b=20 then

OpenStudy (danjs):

the rates of change in area and height are also given, one thing left to solve for

OpenStudy (kayders1997):

Db/dt

OpenStudy (danjs):

any prob like these related rates ones, use some relationship for the variables you need, some may be in terms of others. derivative with respect to what is needed solve

OpenStudy (kayders1997):

I got db/dt is -1.6 m/s?

OpenStudy (danjs):

does that make sense, if the area changes by 2 cm^2/min, and the height increases at 1 cm/min the base has to become smaller?

OpenStudy (kayders1997):

Yes

OpenStudy (danjs):

yeah , it does you can increase time by one min, and see increase area by 2 and h by 1 A = 1/2 b * h 102 = 1/2 * b * 11 b is less than 20

OpenStudy (danjs):

that it, goodluck

OpenStudy (kayders1997):

Thats much easier than trying to think of it from a picture standpoint ill for sure have to write that! So ill remember for the quiz tomorrow

OpenStudy (kayders1997):

Thank you! It's Slowly coming day by day! Thank you for the help! :)